Fermi National Laboratory
 
Operation Summary for April 16, 2012 - April 18, 2012
 

Monday April 16

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), with the SeaQuest experiment in access, and with the NuMI and MiniBooNE experiments taking beam, and with the Meson FTBFT-992 experiment taking beam.

During the day shift, at 8:49 AM, SeaQuest came out of access. A power glitch at 9:34 AM took out the Booster West Gallery Anode and the Booster RF system; everything reset. At 10:18 AM, a NuMI power supply (HV101) tripped off and it took Operators several tries before it reset. At 12:01 PM, Operations received a new Operating Note for the Muon Ring allowing them to take higher intensity 8 GeV beam to their target. A MI person began a g-2 (g minus 2) study at 1:18 PM. The Muon Ring study ended at 3:42 PM. The MiniBooNE beam permit tripped off at 3:49 PM due to an air flow issue (E:12AAIR); it reset. T-992 made an access during this shift.

During the evening shift, at 5:17 PM, Operations reported beam falling out of MI. A MI RF station (MIRF2) tripped off; a RF technician began repairs. At 5:40 PM, Operators rebooted the MI low level (LL) RF system and then rebooted the dampers. Operations tried to run MI beam at low intensity at 6:39 PM and the beam could be seen on the last toroid (TOR852) in the MI-8 transfer line and the MI injection toroid (TOR103), but not beyond. MI personnel managed to get the beam to circulate about 20 turns and reported finding an obstruction in the 602 area. Operators prepared MI for access at 7:40 PM. Operators also prepared SeaQuest for access. MI personnel confirmed that two ferrite rods in a microwave absorber had broken free and one had dropped into the beam. The rods had been cleared and the beamline buttoned back up by 10:47 PM. At 11 PM, Operations reported a problem with BRF2 and called in a RF person to investigate. Vacuum personnel reported at 11:45 PM that they had completed their vacuum check of the MI 602 area and it passed. At 11:46 PM, Operations reported that some ion pumps in MI-10 had tripped off.

Tuesday April 17

The midnight shift began with no beam due to the MI microwave absorber work.

During the midnight shift, Vacuum personnel reported at 2:20 AM that the MI vacuum looked good and they were coming out of the tunnel. Operations reestablished a low intensity MI beam at 3:24 AM; the intensity was increased slowly to prevent the vacuum from degrading. T-1005 began taking beam around 3:25 AM. At 4:40 AM, Operations reported an expected MI vacuum burst in the MI-53 region. Operations established beam to MiniBooNE, NuMI, and T-992 around 5:03 AM. Operations reported at 7:05 AM that MI had resumed sending 100% of its requested beam intensity.

During the day shift, Booster and MI personnel began a phase lock study at 9:49 AM. NuMI began conducting target scans at 2:37 PM; they completed their scans at 3:22 PM. T-992 and SeaQuest made accesses during this shift.

During the evening shift, Operations established beam to T-1005 at 4:43 PM. Operations established beam to MTA at 5:15 PM. At 5:34 PM, LRF4 tripped off; Linac personnel investigated and changed LRF4’s interlock module and its waveform generator. Linac resumed sending beam at 6:12 PM. Operations disabled beam to MTA at 7:58 PM.

Wednesday April 18

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest and T-1005.

During the midnight shift, Operations completed a 750 KeV area line tune at 4:32 AM, but only kept the Linac steer changes in place. Operations disabled beam to T-1005 at 6:46 AM; this ended the Fermi run for both FTBF MTest experiments T-1005 and T-992.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/18/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory